A beam of x rays with wavelength 0.0500 nm isCompton-scattered by the electrons in a sample.

a) At what angle from the incident beam should you look to find x rays with a wavelength of 0.0542 nm?

b) At what angle from the incident beam should you look to find x rays with a wavelength of 0.0521 nm?

c) At what angle from the incident beam should you look to find x rays with a wavelength of 0.0500 nm?
Answer #1

here,

wavelength of the beam , lamdai = 0.05 nm

(a)

lamdaf = 0.0542 nm

let the angle be theta

lamdaf - lamdai = (1-cos(theta)) *h/m0 *c

( 0.0542 - 0.05 ) * 10^-9 = ( 1 -cos(theta)) * 6.626 * 10^-34 /( 9.1*10^-31 * 3*10^8)

theta = 136.89 degree

the angle is 136.89 degree

(b)

lamdaf = 0.0521 nm

let the angle be theta

lamdaf - lamdai = (1-cos(theta)) *h/m0 *c

( 0.0521 - 0.05 ) * 10^-9 = ( 1 -cos(theta)) * 6.626 * 10^-34 /( 9.1*10^-31 * 3*10^8)

theta = 83.68 degree

the angle is 83.68 degree

(c)

lamdaf = 0.05 nm

let the angle be theta

lamdaf - lamdai = (1-cos(theta)) *h/m0 *c

( 0.05 - 0.05 ) * 10^-9 = ( 1 -cos(theta)) * 6.626 * 10^-34 /( 9.1*10^-31 * 3*10^8)

theta = 0 degree

the angle is 0 degree